# TAIA: Large Language Models are Out-of-Distribution Data Learners

### v1

@@ Abstract
Fine-tuning on task-specific question-answer pairs is a predominant method for enhancing the performance of instruction-tuned large language models (LLMs) on downstream tasks. However, in certain specialized domains, such as healthcare or harmless content generation, it is nearly impossible to obtain a large volume of high-quality data that matches the downstream distribution. To improve the performance of LLMs in data-scarce domains with domain-mismatched data, we re-evaluated the Transformer architecture and discovered that not all parameter updates during fine-tuning contribute positively to downstream performance. Our analysis reveals that within the self-attention and feed-forward networks, only the fine-tuned attention parameters are particularly beneficial when the training set’s distribution does not fully align with the test set. Based on this insight, we propose an effective inference-time intervention method: Training All parameters but Inferring with only Attention (TAIA). We empirically validate TAIA using two general instruction-tuning datasets and evaluate it on seven downstream tasks involving math, reasoning, and knowledge understanding across LLMs of different parameter sizes and fine-tuning techniques. Our comprehensive experiments demonstrate that TAIA achieves superior improvements compared to both the fully fine-tuned model and the base model in most scenarios, with significant performance gains. The high tolerance of TAIA to data mismatches makes it resistant to jailbreaking tuning and enhances specialized tasks using general data.
@@ 1 Introduction
Large language models (LLMs) have revolutionized Natural Language Processing (NLP), which has been pretrained on a massive textual corpus and encoded massive world knowledge [8, 1]. These models achieve remarkable zero-shot and few-shot performance across a wide range of tasks [6, 2, 45, 65, 66]. The innovation of instruction tuning, also known as supervised fine-tuning (SFT), has further enhanced the instruction-following capabilities of LLMs [46, 11], simplifying human-LLM interactions. Despite the availability of high-quality data for SFT being limited [7, 86], expanding SFT datasets remains a straightforward method to adapt LLMs for specific tasks [14]. Various SFT datasets, such as Alpaca [49] and Natural Instructions [41, 69], have been manually curated or artificially generated to create more generalized instruction-tuned LLMs.
@@
To address this, we propose avoiding such data dependency by leveraging the intrinsic properties of fine-tuning and developing an inference-time method that does not rely on high-quality in-distribution data. We first conduct an in-depth investigation of the internal Transformer architecture. We find that during fine-tuning, LLMs enhance their instruction-following ability, primarily controlled by the self-attention module [75]. Conversely, most parameterized knowledge is encoded in the feed-forward network (FFN) module [18, 40] during pretraining [56]. Fine-tuning primarily elicits this pretrained knowledge [71, 46, 59], which remains relatively fixed [86]. This insight prompts us to discard the FFN updates during fine-tuning, as only a small portion positively contributes to downstream performance, while most disrupt the knowledge when fine-tuned on task-mismatched data.
A naive approach is to fine-tune only the attention parameters, but this fails to generalize to OOD data due to insufficient exploration of non-linearity. To ensure sufficient learning of non-linearity, we introduce additional FFN parameters during fine-tuning but retain only the beneficial self-attention updates. This strategy, named Training-All-Inferring-only-Attention (TAIA  [^1] Code is available in https://github.com/pixas/TAIA_LLM), achieves both OOD generalization and sufficient optimization space. The comparisons between the proposed method and the vanilla fine-tuning method are shown in Figure 2
@@
Figure 2: Comparison between different fine-tuning and inference methods. Parameters colored with green and yellow represent models finetuned with in-domain and out-of-distribution data, respectively. “ID” and “OOD” represents in-distribution and out-of-distribution, respectively.
@@ 2 Preliminaries
where $W_{g}^{l},W_{u}^{l}\in\mathbb{R}^{d^{\prime}\times d}$, $W_{d}^{l}\in\mathbb{R}^{d\times d^{\prime}}$ and $d^{\prime}$ is the hidden dimension. $\odot$ is the element-wise multiplication and $\mathrm{SiLU}(x)=x\odot\mathrm{sigmoid}(x)$. The feed-forward network uses inverse-bottleneck parameterization methods, which inherently enlarges the representation space and eases the storage of diverse knowledge from different tasks [18, 40].
@@ 3.2 Parameter Selection for Out-of-Distribution (OOD) SFT
Prior research has demonstrated that LLMs possess a wide range of task knowledge [50, 6, 56] after semi-supervised learning on web data. To enhance their proficiency in specific tasks, domain-related instruction-tuning [85] is utilized to improve the knowledge access process [71, 46, 59]. Moreover, studies [75] have shown that self-attention improves LLMs ability to follow instructions through fine-tuning, which aids in effective knowledge elicitation. However, when trained on OOD data, the optimization objective (Eq. 3) involves significant distribution shifts in certain parameter groups. This can disrupt the pre-trained knowledge stored in the Transformer’s feed-forward network (FFN) [5]. Therefore, a balanced approach is to disregard the parameters that are noisily disrupted, while preserving the parameters that contribute to held-out tasks. This approach is already endorsed by existing research [80]. Since fine-tuning prioritizes effective instruction following over absorbing potentially misleading knowledge, it can be inferred that the knowledge acquired by the FFNs during fine-tuning could be considered somewhat redundant.
@@ E.7 Comparing with OOD Generalization Methods
Infer Mode Qwen 1.8b Qwen 7b LLaMA2 7b LLaMA3 8b
@@ F.3 Full Results on Dataset Scales
Appendix G Case Study

### v2

@@ Abstract
Fine-tuning on task-specific question-answer pairs is a predominant method for enhancing the performance of instruction-tuned large language models (LLMs) on downstream tasks. However, in certain specialized domains, such as healthcare or harmless content generation, it is nearly impossible to obtain a large volume of high-quality data that matches the downstream distribution. To improve the performance of LLMs in data-scarce domains with domain-mismatched data, we re-evaluated the Transformer architecture and discovered that not all parameter updates during fine-tuning contribute positively to downstream performance. Our analysis reveals that within the self-attention and feed-forward networks, only the fine-tuned attention parameters are particularly beneficial when the training set’s distribution does not fully align with the test set. Based on this insight, we propose an effective inference-time intervention method: Training All parameters but Inferring with only Attention (TAIA). We empirically validate TAIA using two general instruction-tuning datasets and evaluate it on seven downstream tasks involving math, reasoning, and knowledge understanding across LLMs of different parameter sizes and fine-tuning techniques. Our comprehensive experiments demonstrate that TAIA achieves superior improvements compared to both the fully fine-tuned model and the base model in most scenarios, with significant performance gains. The high tolerance of TAIA to data mismatches makes it resistant to jailbreaking tuning and enhances specialized tasks using general data. Code is available in https://github.com/pixas/TAIA_LLM.
@@ 1 Introduction
Large language models (LLMs) have revolutionized Natural Language Processing (NLP), where LLMs have been pretrained on a massive textual corpus and encoded massive world knowledge [8, 1]. These models achieve remarkable zero-shot and few-shot performance across a wide range of tasks [6, 2, 45, 65, 66]. The innovation of instruction tuning, also known as supervised fine-tuning (SFT), has further enhanced the instruction-following capabilities of LLMs [46, 11], simplifying human-LLM interactions. Despite the availability of high-quality data for SFT being limited [7, 86], expanding SFT datasets remains a straightforward method to adapt LLMs for specific tasks [14]. Various SFT datasets, such as Alpaca [49] and Natural Instructions [41, 69], have been manually curated or artificially generated to create more generalized instruction-tuned LLMs.
@@
To address this, we propose avoiding such data dependency by leveraging the intrinsic properties of fine-tuning and developing an inference-time method that does not rely on high-quality in-distribution data. We first conduct an in-depth investigation of the internal Transformer architecture. We find that during fine-tuning, LLMs enhance their instruction-following ability, primarily controlled by the self-attention module [75]. Conversely, parameterized knowledge is encoded by the key-value intrinsic of the feed-forward network (FFN) module [18, 40] during pretraining [56]. Fine-tuning primarily elicits this pretrained knowledge [71, 46, 59], which remains relatively fixed [86]. This insight prompts us to discard the FFN updates during fine-tuning, as only a small portion positively contributes to downstream performance, while most disrupt the knowledge when fine-tuned on task-mismatched data.
A naive approach is to fine-tune only the attention parameters, but this fails to generalize to OOD data due to insufficient exploration of non-linearity. To ensure sufficient learning of non-linearity, we introduce additional FFN parameters during fine-tuning but retain only the beneficial self-attention updates. This strategy, named Training-All-Inferring-only-Attention (TAIA), achieves both OOD generalization and sufficient optimization space. The comparisons between the proposed method and the vanilla fine-tuning method are shown in Figure 2
@@
Figure 2: Comparison between different fine-tuning and inference methods. Parameters colored with green and yellow represent models finetuned with in-domain and out-of-distribution data, respectively. “ID” and “OOD” represents in-distribution and out-of-distribution, respectively. When we train in-domain data (colored as green) and out-of-domain data (colored as yellow) and evaluate in in-domain test sets and out-of-domain test sets, respectively (The second row; fine-tuning). The vanilla fine-tuning method can only perform well when trained on ID data and evaluated in ID test sets. Compared to vanilla tuning, TAIA can perform generally well on both types of test sets when given OOD data. As a similar approach that only trains attention, TOA (Train-only-attention) performs badly on both types of evaluation sets as it loses sufficient exploration of optimal parameter groups.
@@ 2 Preliminaries
where $W_{g}^{l},W_{u}^{l}\in\mathbb{R}^{d^{\prime}\times d}$, $W_{d}^{l}\in\mathbb{R}^{d\times d^{\prime}}$ and $d^{\prime}$ is the hidden dimension. $\odot$ is the element-wise multiplication and $\mathrm{SiLU}(x)=x\odot\mathrm{sigmoid}(x)$. The feed-forward network uses inverse-bottleneck parameterization methods, which inherently enlarges the representation space and eases the encoding of diverse knowledge from different tasks [18, 40].
@@ 3.2 Parameter Selection for Out-of-Distribution (OOD) SFT
Prior research has demonstrated that LLMs possess a wide range of task knowledge [50, 6, 56] after semi-supervised learning on web data. To enhance their proficiency in specific tasks, domain-related instruction-tuning [85] is utilized to improve the knowledge access process [71, 46, 59]. Moreover, studies [75] have shown that self-attention improves LLMs ability to follow instructions through fine-tuning, which aids in effective knowledge elicitation. However, when trained on OOD data, the optimization objective (Eq. 3) involves significant distribution shifts in certain parameter groups. This can disrupt the pre-trained knowledge encoded through the Transformer’s feed-forward network (FFN) [5]. Therefore, a balanced approach is to disregard the parameters that are noisily disrupted, while preserving the parameters that contribute to held-out tasks. This approach is already endorsed by existing research [80]. Since fine-tuning prioritizes effective instruction following over absorbing potentially misleading knowledge, it can be inferred that the knowledge acquired by the FFNs during fine-tuning could be considered somewhat redundant.
@@ 3.3 Towards Optimal Parameter Selection Strategy
3.4 Implementation of TAIA
During training, the parameters of FFN and self-attention are updated based on max-likelihood modeling:
${\theta}_{ffn}^{\prime},{\theta}_{attn}^{\prime}=\arg\max_{\{\theta\}}\sum_{i=1}^{N}p(\boldsymbol{y}_{i}|\boldsymbol{X}_{i},{\theta}_{ffn},{\theta}_{attn})$ (4)
where $N$ is the number of training samples and $\boldsymbol{X}_{i},\boldsymbol{y}_{i}$ are query and response sequences sampled from any conversational data, like Alpaca-GPT4 or CoT-Collection, respectively. ${\theta}^{\prime}_{(\cdot)}$ is the updated weight in full fine-tuning or the merged weight of LoRA tuning. After training, TAIA only utilizes the updated attention parameters and reuses the pre-trained FFN parameters to perform inference:
$\boldsymbol{y}=\arg\max_{\boldsymbol{y}}\sum_{j=1}^{K}\log p(y_{j}|\boldsymbol{y}_{j-1},\boldsymbol{X},{\theta}_{ffn},{\theta}_{attn}^{\prime})$ (5)
where $K$ is the generated sequence length and $\boldsymbol{X}$ is the query input to LLMs that shares different distributions with the training data. In this scenario, ${\theta}_{ffn}$ is the original parameter of FFN in pre-trained models and ${\theta}_{attn}^{\prime}$ is the updated parameter groups of self-attention in full fine-tuning (or merged weight of self-attention in LoRA tuning).
@@ 4.3 Quantitative Analysis
Table 3: Ablation experiments on different inference modes under two training corpus. We validate the performance of inference modes by considering both general tasks and domain tasks. Bold indicates the optimal result in each subgroup and underline indicates the suboptimal result. Note that even fine-tuned on out-of-domain data, TAIA still achieves optimal results on specific domain tasks and even surpasses the performance of the base model.
Training Data Model Infer Mode Genereal Task Domain Task Average
CMMLU MMLU CEval MMed ZH MMed EN MATH
Qwen1.5-1.8B 52.68 43.62 55.57 62.03 33.78 4.28 41.99
Medical Collection LoRA Vanilla 39.60 27.47 37.74 57.88 29.69 8.24 33.44
TAIA 54.58 44.47 55.57 64.97 36.45 10.20 44.37
TAIF 47.06 42.05 45.62 58.76 32.05 9.06 39.10
TOA 43.37 29.21 39.90 59.54 30.79 7.90 35.12
TOF 41.18 26.64 39.82 58.17 29.22 8.14 33.86
OpenMath LoRA Vanilla 54.04 39.36 50.67 58.03 33.62 7.60 40.55
TAIA 54.35 43.98 56.32 63.81 35.82 11.68 44.33
TAIF 53.46 43.53 54.85 62.84 36.25 7.64 43.09
TOA 53.37 33.73 50.22 57.88 30.56 7.50 38.88
TOF 52.95 37.46 48.37 55.34 31.66 7.48 38.88
@@ 6 Conclusion
Acknowledgments and Disclosure of Funding
We thank Jiangtao Feng for his valuable suggestions on the paper structure. We also thank the anonymous reviewers for their insightful comments and suggestions. This work is supported by National Key R&D Program of China (No. 2022ZD0162101), National Natural Science Foundation of China (No. 62106140) and STCSM (No. 21511101100, No. 22DZ2229005)
@@ E.7 Comparing with OOD Generalization Methods
Infer Mode Qwen1.5 1.8B Qwen1.5 7B LLaMA2 7B LLaMA3 8B
@@ F.3 Full Results on Dataset Scales
Mixing Schedule Methods OOD (20K) OOD (40K) OOD (60K) OOD (80K)
Uniform Vanilla 58.66 58.35 59.66 57.15
TAIA 64.74 64.59 64.62 64.80
Linear Annealing Vanilla 60.42 58.90 60.95 63.16
TAIA 64.97 64.42 64.94 65.18
Table 8: Data mixture ablation on the OOD data ratio. We compare vanilla LoRA tuning with TAIA on two mixture strategies: uniform mixture and linear annealing mixture. TAIA achieves best performance under both settings.
Methods 1-stage 2-stage
ID (20K) OOD (20k) OOD (40k) OOD (60k) OOD (80k)
Vanilla 62.05 28.66 26.53 30.30 30.18
TAIA 64.07 49.71 50.61 47.43 51.05
TOA 27.99 27.64 28.92 30.65
Table 9: Rather than using the one-stage data mixture training setting, we compare TAIA with the vanilla method with a two-stage paradigm, where the ID data is served as the first stage training data and we explicitly separate OOD data in the second stage.
F.4 More Discussion on ID/OOD Data
We here empirically validate that the mixture of ID and OOD data as the training set of TAIA is a sweet configuration compared to other settings. Following the experiment setting in Figure 1(c), we fine-tune Qwen1.5-1.8B models with LoRA where $r=16,\alpha=32$. We design three types of methods to differentiate the ID and OOD data during the fine-tuning process:
  1. 1.
    Rather than mixing the ID and OOD data evenly in Figure 1(c), we adopt linear annealing to gradually decrease the proportion of ID data from 1 to 0 as training progresses.
  2. 2.
    We fine-tune the Qwen1.5-1.8B with 20k ID data in the first stage and 20-80k OOD data in the second stage. Note that we adopt 2-stage fine-tuning based on TAIA-tuned models for better performance.
  3. 3.
    We follow experiment 2’s setting but we fine-tune the Qwen1.5-1.8B with 20-80k OOD data in the first stage and 20k ID data in the second stage.
Table 8 shows that the linear annealing method can mitigate the noise introduced by OOD data, while TAIA can further enhance the model’s ability to utilize OOD data, thereby achieving higher performance on ID data. Table 9 indicates that the 2-stage fine-tuning of OOD data causes a serious performance drop. The first stage not only fits the model to the ID data distribution but also causes the model to lose its generalization ability, making it more sensitive to the noise introduced by OOD data. As a result, the 2-stage fine-tuning method severely degraded the model’s performance. Table 10 demonstrates that compared to the results of Table 9, the models fine-tuned on the OOD data in the 1-stage retain their generalization ability due to the data diversity of the CoT-Collection. However, the results show that such differentiation of the ID and OOD data still fails to improve performance further in the data mixing scenario. As a result, we use the simple yet effective data mixing setting and adopt one-stage training methodology across the whole paper.
Settings Methods OOD(20k) OOD(40k) OOD(60k) OOD(80k)
1-stage TAIA 62.29 61.85 61.03 61.23
2-stage (w/ ID data) Vanilla 59.34 59.63 59.19 59.25
TAIA 60.68 59.60 60.97 61.44
Table 10: We follow Table 9’s setting but we fine-tune the Qwen1.5-1.8B with 20-80k OOD data in the first stage and 20k ID data in the second stage.
Model MATH Medical CoT
Sim/Acc Sim/Acc Sim/Acc
Vanilla 87.50/50.56 78.97/37.74 77.92/24.67
TAIF 96.39/54.85 84.53/45.62 85.83/26.75
TAIA 96.44/56.32 95.91/55.57 91.92/54.38
Table 11: Similarity/performance comparison with pre-trained models, where ‘Sim’ represents the hidden state similarity with pre-trained model and ‘Acc’ is the downstream performance on C-Eval.
F.5 Success of TAIA: A Similarity Perspective
In this experiment, we measure similarity to the pre-trained model. This would measure whether TAIA is essentially regularizing the model towards the pre-trained model (similar to what L2 or EWC would do, but apparently better). We compute the hidden state similarity after each layer and average them just as Figure 5’s setting. The results are shown in Table 11. The results reflect that TAIA regularizes the fine-tuned model in an implicit manner without disturbing the learning and parameter exploration, which happens in other regularization methods like L2 and EWC.
Table 12: Model size scaling of TAIA. We choose 7B, 14B, 32B of Qwen1.5 series as the pre-trained models. TAIA maintains the best performance based on even larger pre-trained models.
Model Infer Mode Reasoning Knowledge Average
MATH BBH CQA LogiQA SVAMP MMB MMLU
Qwen1.5-7B Base 20.30 30.76 78.30 42.70 54.90 45.09 57.69 47.11
LoRA 17.90 36.09 77.31 37.33 57.10 44.85 54.89 46.50
TAIA 24.98 43.46 77.31 41.78 67.20 46.90 57.29 51.27
Qwen1.5-14B Base 45.98 43.88 77.72 47.16 83.60 51.06 66.05 59.35
LoRA 38.74 41.65 74.45 41.78 82.80 48.15 63.71 55.90
TAIA 55.51 46.06 77.31 46.39 83.10 52.55 65.48 60.92
Qwen1.5-32B Base 41.22 48.78 80.92 50.23 87.60 61.04 73.03 63.26
LoRA 39.12 46.29 77.81 49.31 82.60 59.54 71.02 60.81
TAIA 42.70 52.63 82.31 48.69 86.20 61.98 72.63 63.88
F.6 Model Scaling of TAIA
To investigate what happens with even larger scale models, we conduct experiments based on the 14B and 32B sizes of Qwen1.5 using LoRA tuning and Alpaca-GPT4 data and show the results in Table 12 shows that TAIA still outperforms both the base model and LoRA tuning model with Alpaca-GPT4 data, especially in reasoning tasks like MATH and BBH, which further verifies the effectiveness of TAIA.
@@ F.6 Model Scaling of TAIA
Table 15: Variability of TAIA. TAIA performs generally more superior to the vanilla LoRA fine-tuning.
Setting Infer Mode Reasoning Knowledge Average
MATH BBH CQA LogiQA SVAMP MMB MMLU
Base 20.30 30.76 78.30 42.70 54.90 45.09 57.69 47.11
Run1 LoRA 17.90 36.09 77.31 37.33 57.10 44.85 54.89 46.50
TAIA 24.98 43.46 77.31 41.78 67.20 46.90 57.29 51.27
Run2 LoRA 18.44 38.90 75.35 36.25 56.80 44.62 55.41 46.54
TAIA 26.52 43.33 75.76 41.56 67.60 46.27 57.36 51.20
Run3 LoRA 18.36 38.66 76.41 36.87 56.70 44.46 55.39 46.69
TAIA 27.02 43.56 76.41 41.56 67.90 46.11 57.12 51.38
F.7 Variability of TAIA
In this section, we validate that TAIA is a robust method that is hardly affected by random seeds. To this end, we choose Qwen1.5 7B as the base model and Alpaca-GPT4 as the training data, and repeat the training process for three runs. The results are shown in Table 15. We observe that the improvements do not vary among different runs, which emphasizes the robustness of TAIA.
Appendix G Formalize the utility of TAIA
Suppose an LLM $p_{\theta}$ containing pretrained weight $\theta_{0}$, the vanilla LoRA-tuned model yields $\Delta\theta_{0}$ weight which is to be merged back to pretrained weight and has a relatively small norm. Suppose a simplified neural network layer with nonlinear operators, a layer normalization layer $\mathrm{LayerNorm}(X)=\frac{X-\mu}{\sigma}$ and a residual connection:
$M_{W}(X)=\text{Act}(Wx)$ (10)
As the magnitude of $\Delta\theta_{0}$ is quite small compared with $\theta_{0}$, we perform a first-order Taylor expansion on $M_{\theta_{0}+\Delta\theta_{0}}(X)$:
$M_{\theta_{0}+\Delta\theta_{0}}(X)\approx M_{\theta_{0}}(X)+J_{\theta_{0}}(X)\Delta\theta_{0}$ (11)
where $J_{\theta_{0}}(X)$ is the Jacobian matrix of $M_{\theta_{0}}(X)$ for $\theta_{0}$. We define $z=X+M_{\theta_{0}}(X)$ and $z^{\prime}=X+M_{\theta_{0}}(X)+J_{\theta_{0}}(X)\Delta\theta_{0}$ to compare $\mathrm{LayerNorm(z)}$ and $\text{LayerNorm}(z^{\prime})$. Compare the addition of $\Delta\theta_{0}$ inside the transformer module:
$X=\text{LayerNorm}(z)\quad X^{\prime}=\text{LayerNorm}(z^{\prime})$ (12)
Considering $z^{\prime}=z+J_{\theta_{0}}(X)\Delta\theta_{0}$, we have
$\mu_{z^{\prime}}\approx\mu_{z}+\mu_{J_{\theta_{0}}(X)\Delta\theta_{0}}\quad\sigma_{z^{\prime}}\approx\sigma_{z}$ (13)
Since $|\Delta\theta_{0}|$ is quite small compared to $\theta_{0}$, $J_{\theta_{0}}(X)\Delta\theta_{0}$ is also of small norm, which can be considered to be ignored for the mean and standard deviation. Based on it, we have
$\text{LayerNorm}(z^{\prime})\approx\text{LayerNorm}(z)$
In other words, if the updated parameter is small enough compared to the pretrained model, the output distribution after each submodule remains consistent with the pretrained model. Meanwhile, we need to prove that the perturbation brought by attention is far lower than that brought by FFN so that the removal of FFN results in higher improvement than the removal of attention modules. Specifically, omitting the multi-head mechanism, self-attention is formalized as such:
$\text{Attn}(X)=\text{SoftMax}\left(\frac{XW_{q}(XW_{k})^{\top}}{\sqrt{d_{k}}}\right)XW_{v}$ (14)
where $X\in\mathbb{R}^{n\times d}$, and $W_{q},W_{k}\in\mathbb{R}^{d\times d_{k}},W_{v}\in\mathbb{R}^{d\times d_{v}}$, respectively.
For small perturbations $\Delta W_{q},\Delta W_{k},\Delta W_{v}$, the perturbed Attention output is:
$\text{Attn}_{\Delta}(X)=\text{SoftMax}\left(\frac{X(W_{q}+\Delta W_{q})(X(W_{k}+\Delta W_{k}))^{\top}}{\sqrt{d_{k}}}\right)X(W_{v}+\Delta W_{v})$ (15)
Using the first-order Taylor expansion, we have:
$\Delta\text{Attn}(X)\approx\frac{\partial\text{Attn}(X)}{\partial W_{q}}\Delta W_{q}+\frac{\partial\text{Attn}(X)}{\partial W_{k}}\Delta W_{k}+\frac{\partial\text{Attn}(X)}{\partial W_{v}}\Delta W_{v}$ (16)
Using the first-order Taylor expansion and the partial derivatives of softmax, we obtain the perturbation bound of attention:
$\|\Delta\text{Attn}(X)\|\leq\frac{1}{\sqrt{d_{k}}}(\|(\text{diag}(P)-PP^{\top})X(XW_{k})^{\top}XW_{v}\|\|\Delta W_{q}\|$
$+\|(\text{diag}(P)-PP^{\top})XW_{q}^{\top}XXW_{v}\|\|\Delta W_{k}\|)+\|P^{\top}X\|\|\Delta W_{v}\|$
where $P=\text{SoftMax}(\frac{XW_{q}(XW_{k})^{\top}}{\sqrt{d_{k}}})$ We simplify the bound as such $\|\Delta\text{Attn}(X)\|\leq C_{attn}(\|\Delta W_{q}\|+\|\Delta W_{k}\|+\|\Delta W_{v}\|)$ where $C_{attn}=\frac{\|X\|^{3}\|W_{k}\|\|W_{v}\|}{\sqrt{d_{k}}}\|\text{diag}(P)-PP^{\top}\|$.
In terms of FFN modules, we use ReLU as the activation function instead of SwiGLU in FFN, which is defined as:
$\text{FFN}(X)=\text{ReLU}(W_{1}(\text{ReLU}(W_{2}x+b_{2}))+b_{1})$ (17)
where $W_{1}\in\mathbb{R}^{d\times 4d}$, $W_{2}\in\mathbb{R}^{4d\times d}$. For small perturbations $\Delta W_{1},\Delta W_{2},\Delta b_{1},\Delta b_{2}$, the perturbed FFN output is:
$\text{FFN}_{\Delta}(X)=\text{ReLU}((W_{1}+\Delta W_{1})(\text{ReLU}((W_{2}+\Delta W_{2})X+b_{2}+\Delta b_{2}))+b_{1}+\Delta b_{1})$ (18)
Similar to the above analysis, we obtain the perturbation bound of FFN modules:
$\|\Delta\text{FFN}(X)\|\leq\|f^{\prime}(Y)XW_{1}f^{\prime}(XW_{2}+b_{2})\|\|\Delta W_{2}\|+\|f^{\prime}(Y)W_{1}f^{\prime}(X_{2}+b_{2})\|\|\Delta b_{2}\|$
$+\|f^{\prime}(Y)Y\|\|\Delta W_{1}\|+\|f^{\prime}(Y)\|\|\Delta b_{1}\|$
where $f^{\prime}$ is the derivative of ReLU and $Y=\text{ReLU}(XW_{2}+b_{2})$ and we simplify the bound as such: $\|\Delta\text{FFN}(X)\|\leq C_{ffn}(\|\Delta W_{2}\|+\|\Delta W_{1}\|+\|\Delta b_{1}\|+\|\Delta b_{2}\|)$ where $C_{ffn}=nd\|X\|\cdot\|W_{1}|b_{1}\|\cdot\|W_{2}|b_{2}\|$.
The SoftMax function’s normalization limits the perturbation’s amplification and hence limits the $\|\text{diag}(P)-PP^{\top}\|\leq 1$ and the overall perturbation. Besides, $\|X\|_{2}^{2}\sim\chi^{2},\|X\|\ \approx\sqrt{d}$. Due to ReLU’s non-linear activation, perturbations are significantly amplified around the non-linear activations, especially in the high-dimensional space. Based on the above comparison
$C_{attn}\approx d\|W_{k}\|\|W_{v}\|\|\text{diag}(P)-PP^{\top}\|\ll C_{ffn}=nd\sqrt{d}\cdot\|W_{1}|b_{1}\|\cdot\|W_{2}|b_{2}\|$ (19)
which implies
$\|\Delta\text{Attn}(X)\|<\|\Delta\text{FFN}(X)\|$ (20)
Thus, for equal-magnitude parameter perturbations, the output perturbation bound of the Attention module is less than that of the FFN module. This indicates that the FFN updates will introduce great perturbation to the distribution of transformers and narrow the similarity between tuned models and pre-trained models.
Appendix H Case Study